OpenStudy (anonymous):

write the equation of the line thorugh (0,1) and perpendicular to 2x-3y=5

myininaya (myininaya):

write 2x-3y=5 in y=mx+b form so we can determine the slope of this line determining the slope of that line will help us to determine the slope of the line perpendicular to it

OpenStudy (amistre64):

perp; eh....get the slope here and then turn it over and negate it.. 2/3...need -3/2 for the other line 3x+2y = 2

OpenStudy (amistre64):

the slope is -a/b ax +by = c

OpenStudy (anonymous):

you guys are life savers!!!!!

myininaya (myininaya):

but we want the line that is running through (0,1) y=-3x/2+b plug x=0 and y=1 to determine y intercept